OpenStudy (anonymous):

Find the distance between J and K. A. 13.42 B. 8.49 C. 7.21 D. 12.65

OpenStudy (anonymous):

do you have an image?

OpenStudy (anonymous):

there

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

tell me if you get it correect

OpenStudy (anonymous):

it was wrong but its okay (:

OpenStudy (anonymous):

It was B huh? :(

OpenStudy (anonymous):

yeah (:

OpenStudy (anonymous):

Yea i thought it was be but i went with C sorry :(

OpenStudy (anonymous):

oh no its okay (: i still got an 86.7 percent on my test

OpenStudy (anonymous):

Oh thats good anyways, what grade you in?

OpenStudy (radar):

You are asked to find the distance between 2 points that are located at (5,9) and (-1, 3). Note that is also the hypotenuse of a right triangle whose legs are each 6 units long. Use the pythagorean theorem

OpenStudy (radar):

\[\sqrt{6^{2}+6^{2}}\]

OpenStudy (radar):

\[\sqrt{72}=6\sqrt{2}\]